# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ Abstract.
Using Givental’s non-linear Maslov index we define a sequence of spectral selectors on the universal cover of the identity component of the contactomorphism group of any lens space. As applications, we prove that the standard Reeb flow is a geodesic for the discriminant and oscillation norms, and we define a stably unbounded conjugation invariant spectral pseudonorm.
@@ 1. Introduction
Since the weights $\underline{w}$ do not play a particular role in the discussion, we denote $L_{k}^{2n-1}(\underline{w})$ simply by $L_{k}^{2n-1}$. We endow $L_{k}^{2n-1}$ with its canonical contact structure $\xi_{0}$, the kernel of the contact form $\alpha_{0}$ whose pullback $\bar{\alpha}_{0}$ by the projection $\mathbb{S}^{2n-1}\rightarrow L_{k}^{2n-1}$ is equal to the pullback of $\sum_{j=1}^{n}x_{j}dy_{j}-y_{j}dx_{j}$ by the inclusion $\mathbb{S}^{2n-1}\hookrightarrow\mathbb{R}^{2n}$. We denote by $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ the universal cover of the identity component $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ of the contactomorphism group. The non-linear Maslov index is a quasimorphism
$\mu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathbb{Z}\,,$
@@
$c_{j}:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\to\mathbb{R}$
that associate to every element of $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ a real number belonging to its action spectrum. Recall that the action spectrum of a contactomorphism $\phi$ of a contact manifold $(M,\xi)$ with respect to a contact form $\alpha$ is the set $\mathcal{A}_{\alpha}(\phi)$ of real numbers $T$ that are translations of translated points of $\phi$ with respect to $\alpha$. We denote by
@@
$\mathcal{L}:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\widetilde{\operatorname{Cont}_{0}}\big{(}\mathbb{S}^{2n-1},\bar{\xi}_{0}=\ker(\bar{\alpha}_{0})\big{)}$
be the map that sends $\widetilde{\phi}=[\{\phi_{t}\}_{t\in[0,1]}]$ to the element of $\widetilde{\operatorname{Cont}_{0}}(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ represented by the lift of $\{\phi_{t}\}_{t\in[0,1]}$ to $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$. For $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ we denote
@@
Since the Reeb flow of $\alpha_{0}$ on $L_{k}^{2n-1}$ is periodic of period $\frac{2\pi}{k}$ and the Reeb flow of $\bar{\alpha}_{0}$ on $\mathbb{S}^{2n-1}$ is periodic of period $2\pi$, $\mathcal{A}(\widetilde{\phi})$ and $\bar{\mathcal{A}}(\widetilde{\phi})$ are invariant by translation by $\frac{2\pi}{k}$ and $2\pi$ respectively. For a real number $T$ we denote
$\big{\lceil}T\big{\rceil}_{\frac{2\pi}{k}}=\frac{2\pi}{k}\left\lceil\frac{k}{2\pi}\,T\right\rceil\quad\text{ and }\quad\big{\lfloor}T\big{\rfloor}_{\frac{2\pi}{k}}=\frac{2\pi}{k}\left\lfloor\frac{k}{2\pi}\,T\right\rfloor\,,$
thus $\lceil T\rceil_{\frac{2\pi}{k}}$ and $\lfloor T\rfloor_{\frac{2\pi}{k}}$ are respectively the smallest multiple of $\frac{2\pi}{k}$ greater or equal than $T$ and the greatest multiple of $\frac{2\pi}{k}$ smaller or equal than $T$.
Before stating our main result we recall that, since $(L_{k}^{2n-1},\xi_{0})$ is orderable, the relation $\leq$ on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ defined by posing $\widetilde{\phi}\leq\widetilde{\psi}$ if there is a non-negative contact isotopy representing $\widetilde{\psi}\,\cdot\,\widetilde{\phi}^{-1}$ is a bi-invariant partial order; this is the partial order that appears in point (vii) below. Recall also that a translated point $p$ of a contactomorphism $\phi$ of a contact manifold $(M,\xi)$ with respect to a contact form $\alpha$ is said to be non-degenerate for a translation $T$ if there is no vector $X\in T_{p}M\smallsetminus\{0\}$ such that $(r^{\alpha}_{-T}\circ\phi)_{\ast}(X)=X$ and $dg(X)=0$, where $g$ is the conformal factor of $\phi$, i.e. the function defined by the relation $\phi^{\ast}\alpha=e^{g}\alpha$. In the case of $(L_{k}^{2n-1},\xi_{0})$ or $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$, if a translated point of a contactomorphism with respect to $\alpha_{0}$ or $\bar{\alpha}_{0}$ is non-degenerate for a certain translation then it is non-degenerate for all the translations; we then just say that it is non-degenerate. For any $T\in\mathbb{R}$ we denote
@@
where $\{r_{t}\}$ is the Reeb flow on $L_{k}^{2n-1}$ with respect to $\alpha_{0}$. Moreover, we denote by $\widetilde{\mathrm{id}}$ the identity on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$.
@@
$\big{\{}\,c_{j}:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\to\mathbb{R}\,,\;j\in\mathbb{Z}\,\big{\}}$
@@
    $c_{-2n+1}(\widetilde{\mathrm{id}})=c_{0}(\widetilde{\mathrm{id}})=0\,.$
@@
    Relation with translated points: if all the translated points of $\Pi\big{(}\mathcal{L}(\widetilde{\phi})\big{)}$ are non-degenerate then the spectral selectors $\{\,c_{j}(\widetilde{\phi})\,,j\in\mathbb{Z}\,\}$ are all distinct. On the other hand, if
@@
    for some $j$ and $11\leq m\leq 2n-$ 1 and either $k$ is even or $j$ is odd or $m>1$ then $\Pi\big{(}\mathcal{L}(\widetilde{\phi})\big{)}$ has infinitely many translated points of translation $T$.
@@
    Continuity: if $\widetilde{\phi}\cdot\widetilde{\psi}^{-1}$ is represented by a contact isotopy with Hamiltonian function $H_{t}:L_{k}^{2n-1}\rightarrow\mathbb{R}$ with respect to $\alpha_{0}$ then
@@
    $c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\leq c_{j}\big{(}\widetilde{\phi}\big{)}+\left\lceil c_{l}\big{(}\widetilde{\psi}\big{)}\right\rceil_{\frac{2\pi}{k}}\,,$
@@
    $\left\lceil c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\right\rceil_{\frac{2\pi}{k}}\leq\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{\frac{2\pi}{k}}+\left\lceil c_{l}\big{(}\widetilde{\psi}\big{)}\right\rceil_{\frac{2\pi}{k}}\,.$
@@
    $\left\lceil c_{j}\big{(}\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1}\big{)}\right\rceil_{\frac{2\pi}{k}}=\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{\frac{2\pi}{k}}\,.$
@@
    $\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{\frac{2\pi}{k}}=-\left\lfloor c_{-j-(2n-1)}\big{(}\widetilde{\phi}^{-1}\big{)}\right\rfloor_{\frac{2\pi}{k}}\,.$
@@
satisfying properties analogue to those of Theorem 1.1, with stronger statements for (ix), (x) and (xi) not involving the $\frac{2\pi}{k}$-floors and ceilings. Such spectral invariants coincide with the ones defined by the first author in [2], and their projections to $\mathbb{S}^{1}$ coincide with the rotation numbers defined by Théret in [22]. Moreover, their properties are analogue to those satisfied by the spectral invariants defined with Floer homology by Entov and Polterovich in [11]. The fact that in the contact case the statements of the triangle inequality, conjugation invariance and Poincaré duality properties are weaker than in the symplectic case and involve the $\frac{2\pi}{k}$-floors and ceilings is similar to what happens for the spectral selectors of compactly supported contactomorphisms of $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ defined by the third author in [19]. Indeed, these spectral selectors are contact analogues of the spectral selectors of compactly supported Hamiltonian symplectomorphisms of $(\mathbb{R}^{2n},\omega_{0})$ defined by Viterbo in [24], but they satisfy weaker versions of the triangle inequality, conjugation invariance and Poincaré duality properties involving their (integral) floors and ceilings. Roughly speaking, this can be explained as follows. In both cases the contact spectral selectors are generalizations of the symplectic ones, in the sense that the symplectic spectral selectors of Hamiltonian isotopies of $(\mathbb{CP}^{n-1},\omega_{0})$ and $(\mathbb{R}^{2n},\omega_{0})$ coincide with the contact spectral selectors of their lifts to $(L_{k}^{2n-1},\xi_{0})$ and $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ respectively. The fact that the contact spectral selectors satisfy weaker versions of the triangle inequality, conjugation invariance and Poincaré duality properties involving their floors and ceilings with respect to the period of the Reeb flow is due to the fact that, while the lifts of Hamiltonian isotopies of $(\mathbb{CP}^{n-1},\omega_{0})$ and $(\mathbb{R}^{2n},\omega_{0})$ are exactly the contact isotopies that commute with the standard Reeb flows, general contact isotopies commute with the Reeb flow at time $t$ only when $t$ is a multiple of the period of the Reeb flow. For conjugation invariance, for instance, while in the symplectic case the action spectrum is invariant by conjugation, in the contact case this is in general not true: the translated points of a contactomorphism are in general not in bijection with those of a conjugation. However, if the Reeb flow is periodic then the translated points of translation equal to the period of the Reeb flow are discriminant points, which are invariant by conjugation, and this fact can be used to prove that the corresponding floor and ceiling of the spectral selectors are invariant by conjugation (see also the discussion in [19]).
@@
Properties (i), (iii) and (vi) imply that every contactomorphism of $(L_{k}^{2n-1},\xi_{0})$ contact isotopic to the identity has at least $n$ translated points with respect to $\alpha_{0}$, and at least $2n$ if $k$ is even or if all the translated points are non-degenerate. We thus recover the corresponding result of [16], but not the optimal bound obtained by the first author in [3], where it is proved that every contactomorphism of $(L_{k}^{2n-1},\xi_{0})$ contact isotopic to the identity has at least $2n$ translated points.
@@
Suppose that $k$ is prime. Recall that the cohomological index $\operatorname{ind}(A)$ of a subset $A$ of $L_{k}^{2n-1}$ is the dimension over $\mathbb{Z}_{k}$ of the image of the map $\check{H}^{\ast}(L_{k}^{2n-1};\mathbb{Z}_{k})\rightarrow\check{H}^{\ast}(A;\mathbb{Z}_{k})$ on Čech cohomology induced by the inclusion $A\hookrightarrow L_{k}^{2n-1}$. As we will see, property (iii) can be refined in this case as follows: if
@@
As a first application of Theorem 1.1 we prove that the standard Reeb flow on $(L_{k}^{2n-1},\xi_{0})$ is a geodesic for the discriminant and oscillation norms introduced in [9]. The definition of the discriminant norm $\nu_{\operatorname{dis}}$ and of the oscillation pseudonorm $\nu_{\operatorname{osc}}$ on the universal cover of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ are recalled in Section 4 below, as well as the definition of the discriminant and oscillation lengths of contact isotopies. Recall also from [9, Proposition 3.2] that the oscillation pseudonorm is non-degenerate if and only if $(M,\xi)$ is orderable; in particular, it is thus a norm for lens spaces. As in [5], we say that a contact isotopy of a closed orderable contact manifold is a geodesic for the discriminant or for the oscillation norm if its discriminant or oscillation length is equal to the discriminant or oscillation norm of the element of the universal cover it represents. In other words, a contact isotopy is a geodesic for the discriminant or oscillation norm if it minimizes the discriminant or oscillation length in its homotopy class with fixed endpoints. In [5] it is proved that certain contact isotopies of $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ are geodesics for the discriminant and oscillation norms. We obtain here a similar result for lens spaces, answering a question in [9].
@@
In [9] and [16] respectively it is proved that the discriminant and oscillation norms on real projective spaces and on general lens spaces are unbounded, by showing that the classes in the universal cover represented by higher iterations of the Reeb flow have bigger and bigger discriminant and oscillation norms. More precisely, it is proved in [16] that for every $N$ the discriminant and oscillation norms on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ of $\widetilde{r_{6\pi N}}$ and $\widetilde{r_{20\pi N}}$ respectively are at least equal to $N+1$. Since the discriminant length of $\{r_{6\pi Nt}\}_{t\in[0,1]}$ is $13Nk+$ 1 and the oscillation length of $\{r_{20\pi Nt}\}_{t\in[0,1]}$ is $110Nk+$ 1, the results in [16] (as well as the previous ones in [9]) left open the question of whether there exist contact isotopies in the same homotopy class with fixed endpoints as $\{r_{6\pi Nt}\}_{t\in[0,1]}$ or $\{r_{20\pi Nt}\}_{t\in[0,1]}$ having shorter discriminant or oscillation lengths. Note that this is what happens for the sphere $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$: the $N$-th iteration $\{r_{2\pi Nt}\}_{t\in[0,1]}$ of the Reeb flow $\{r_{2\pi t}\}_{t\in[0,1]}$ of $\bar{\alpha}_{0}$ has discriminant and oscillation length $N+1$, but by [9, Proposition 4.3] the discriminant norm and the oscillation pseudonorm of [$[\{r_{2\pi Nt}\}_{t\in[0,1]}]$ are smaller or equal than $4$; in other words, there exist contact isotopies of $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ in the same homotopy class with fixed endpoints as certain iterations the Reeb flow having strictly shorter discriminant and oscillation length. As an application of Theorem 1.1, in Section 4 we show that for lens spaces this is not possible. More precisely, we show that Theorem 1.1 implies the following result.
@@
For every real number $T$, the Reeb flow $\{r_{Tt}\}_{t\in[0,1]}$ on $(L_{k}^{2n-1},\xi_{0})$ is a geodesic for the discriminant and oscillation norms. In particular,
@@
Using the spectral selectors of Theorem 1.1 we also define a stably unbounded conjugation invariant pseudonorm on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. More precisely, posing $c_{-}=c_{-2n+1}$ and $c_{+}=c_{0}$ we prove the following result.

### v2

@@ Abstract.
Using Givental’s non-linear Maslov index we define a sequence of spectral selectors on the universal cover of the identity component of the contactomorphism group of any lens space. As applications, we prove for lens spaces with equal weights that the standard Reeb flow is a geodesic for the discriminant and oscillation norms, and we define for general lens spaces a stably unbounded conjugation invariant spectral pseudonorm.
@@ 1. Introduction
We endow $L_{k}^{2n-1}(\underline{w})$ with its canonical contact structure $\xi_{0}$, the kernel of the contact form $\alpha_{0}$ whose pullback $\bar{\alpha}_{0}$ by the projection $\mathbb{S}^{2n-1}\rightarrow L_{k}^{2n-1}(\underline{w})$ is equal to the pullback of $\sum_{j=1}^{n}x_{j}dy_{j}-y_{j}dx_{j}$ by the inclusion $\mathbb{S}^{2n-1}\hookrightarrow\mathbb{R}^{2n}$. We denote by $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ the universal cover of the identity component $\operatorname{Cont}_{0}(L_{k}^{2n-1}\big{(}\underline{w}),\xi_{0}\big{)}$ of the contactomorphism group. The non-linear Maslov index is a quasimorphism
$\mu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{Z}\,,$
@@
$c_{j}:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\to\mathbb{R}$
that associate to every element of $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ a real number belonging to its action spectrum. Recall that the action spectrum of a contactomorphism $\phi$ of a contact manifold $(M,\xi)$ with respect to a contact form $\alpha$ is the set $\mathcal{A}_{\alpha}(\phi)$ of real numbers that are translations of translated points of $\phi$ with respect to $\alpha$. We denote by
@@
$\mathcal{L}:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\widetilde{\operatorname{Cont}_{0}}\big{(}\mathbb{S}^{2n-1},\bar{\xi}_{0}=\ker(\bar{\alpha}_{0})\big{)}$
be the map that sends $\widetilde{\phi}=[\{\phi_{t}\}_{t\in[0,1]}]$ to the element of $\widetilde{\operatorname{Cont}_{0}}(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ represented by the lift of $\{\phi_{t}\}_{t\in[0,1]}$ to $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$. For $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ we denote
@@
The sets $\mathcal{A}(\widetilde{\phi})$ and $\bar{\mathcal{A}}(\widetilde{\phi})$ are invariant by translation by $T_{\underline{w}}$ and $2\pi$ respectively, where $T_{\underline{w}}$ denotes the period of the Reeb flow of $\alpha_{0}$ on $L_{k}^{2n-1}(\underline{w})$. For a real number $T$ we denote
$\big{\lceil}T\big{\rceil}_{T_{\underline{w}}}=T_{\underline{w}}\,\left\lceil\,\frac{T}{T_{\underline{w}}}\,\right\rceil\quad\text{ and }\quad\big{\lfloor}T\big{\rfloor}_{T_{\underline{w}}}=T_{\underline{w}}\,\left\lfloor\,\frac{T}{T_{\underline{w}}}\,\right\rfloor\,,$
thus $\lceil T\rceil_{T_{\underline{w}}}$ and $\lfloor T\rfloor_{T_{\underline{w}}}$ are respectively the smallest multiple of $T_{\underline{w}}$ greater or equal than $T$ and the greatest multiple of $T_{\underline{w}}$ smaller or equal than $T$.
Before stating our main result we recall that, since $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is orderable, the relation $\leq$ on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ defined by posing $\widetilde{\phi}\leq\widetilde{\psi}$ if there is a non-negative contact isotopy representing $\widetilde{\psi}\,\cdot\,\widetilde{\phi}^{-1}$ is a bi-invariant partial order; this is the partial order that appears in point (vii) below. Recall also that a translated point $p$ of a contactomorphism $\phi$ of a contact manifold $(M,\xi)$ with respect to a contact form $\alpha$ is said to be non-degenerate for a translation $T$ if there is no vector $X\in T_{p}M\smallsetminus\{0\}$ such that $(r^{\alpha}_{-T}\circ\phi)_{\ast}(X)=X$ and $dg(X)=0$, where $g$ is the conformal factor of $\phi$, i.e. the function defined by the relation $\phi^{\ast}\alpha=e^{g}\alpha$. In the case of $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$, if a translated point of a contactomorphism with respect to $\bar{\alpha}_{0}$ is non-degenerate for a certain translation then it is non-degenerate for all the translations; we then just say that it is non-degenerate. For any $T\in\mathbb{R}$ we denote
@@
where $\{r_{t}\}$ is the Reeb flow on $L_{k}^{2n-1}(\underline{w})$ of $\alpha_{0}$. Moreover, we denote by $\widetilde{\mathrm{id}}$ the identity on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$.
@@
$\big{\{}\,c_{j}:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\to\mathbb{R}\,,\;j\in\mathbb{Z}\,\big{\}}$
@@
    $c_{0}(\widetilde{\mathrm{id}})=0\,.$
@@
    Relation with translated points: if all the translated points of $\Pi\big{(}\mathcal{L}(\widetilde{\phi})\big{)}$ with respect to $\bar{\alpha}_{0}$ are non-degenerate then the spectral selectors $\{\,c_{j}(\widetilde{\phi})\,,j\in\mathbb{Z}\,\}$ are all distinct. On the other hand, if
@@
    for some $j$ and $11\leq m\leq 2n-$ 1 and either $k$ is even or $j$ is odd or $m>1$ then $\Pi\big{(}\mathcal{L}(\widetilde{\phi})\big{)}$ has infinitely many translated points of translation $T$ with respect to $\bar{\alpha}_{0}$.
@@
    Continuity: if $\widetilde{\phi}\cdot\widetilde{\psi}^{-1}$ is represented by a contact isotopy with Hamiltonian function $H_{t}:L_{k}^{2n-1}(\underline{w})\rightarrow\mathbb{R}$ with respect to $\alpha_{0}$ then
@@
    $c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\leq c_{j}\big{(}\widetilde{\phi}\big{)}+\left\lceil c_{l}\big{(}\widetilde{\psi}\big{)}\right\rceil_{T_{\underline{w}}}\,,$
@@
    $\left\lceil c_{j+l}\big{(}\widetilde{\phi}\cdot\widetilde{\psi}\big{)}\right\rceil_{T_{\underline{w}}}\leq\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{T_{\underline{w}}}+\left\lceil c_{l}\big{(}\widetilde{\psi}\big{)}\right\rceil_{T_{\underline{w}}}\,.$
@@
    $\left\lceil c_{j}\big{(}\widetilde{\psi}\cdot\widetilde{\phi}\cdot\widetilde{\psi}^{-1}\big{)}\right\rceil_{T_{\underline{w}}}=\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{T_{\underline{w}}}\,.$
@@
    $\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{T_{\underline{w}}}=-\left\lfloor c_{-j-(2n-1)}\big{(}\widetilde{\phi}^{-1}\big{)}\right\rfloor_{T_{\underline{w}}}\,.$
@@
satisfying properties analogue to those of Theorem 1.1, with stronger statements for (ix), (x) and (xi) not involving the $T_{\underline{w}}$-floors and ceilings. Such spectral invariants coincide with the ones defined by the first author in [2], and their projections to $\mathbb{S}^{1}$ coincide with the rotation numbers defined by Théret in [23]. Moreover, their properties are analogue to those satisfied by the spectral invariants defined with Floer homology by Entov and Polterovich in [12]. The fact that in the contact case the statements of the triangle inequality, conjugation invariance and Poincaré duality properties are weaker than in the symplectic case and involve the $T_{\underline{w}}$-floors and ceilings is similar to what happens for the spectral selectors of compactly supported contactomorphisms of $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ defined by the third author in [20]. Indeed, these spectral selectors are contact analogues of the spectral selectors of compactly supported Hamiltonian symplectomorphisms of $(\mathbb{R}^{2n},\omega_{0})$ defined by Viterbo in [25], but they satisfy weaker versions of the triangle inequality, conjugation invariance and Poincaré duality properties involving their (integral) floors and ceilings. Roughly speaking, this difference with respect to the symplectic case can be explained as follows. If we see $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ and $\big{(}L_{k}^{2n-1}(1,\cdots,1),\xi_{0}\big{)}$ as prequantizations of $(\mathbb{R}^{2n},\omega_{0})$ and $(\mathbb{CP}^{n-1},\omega_{0})$ respectively then in both cases the contact spectral selectors are generalizations of the symplectic ones, in the sense that the symplectic spectral selectors of Hamiltonian isotopies of $(\mathbb{R}^{2n},\omega_{0})$ and $(\mathbb{CP}^{n-1},\omega_{0})$ coincide with the contact spectral selectors of their lifts to $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ and $\big{(}L_{k}^{2n-1}(1,\cdots,1),\xi_{0}\big{)}$ respectively. The fact that the contact spectral selectors satisfy weaker versions of the triangle inequality, conjugation invariance and Poincaré duality properties involving their floors and ceilings with respect to the period of the Reeb flow is due to the fact that, while the lifts of Hamiltonian isotopies of $(\mathbb{R}^{2n},\omega_{0})$ and $(\mathbb{CP}^{n-1},\omega_{0})$ are exactly the contact isotopies that commute with the standard Reeb flows, general contact isotopies commute with the Reeb flow at time $t$ only when $t$ is a multiple of the period of the Reeb flow. For conjugation invariance, for instance, while in the symplectic case the action spectrum is invariant by conjugation, in the contact case this is in general not true: the translated points of a contactomorphism are in general not in bijection with those of a conjugation. However, if the Reeb flow is periodic then the translated points of translation equal to the period of the Reeb flow are discriminant points, which are invariant by conjugation, and this fact can be used to prove that the corresponding floor and ceiling of the spectral selectors are invariant by conjugation (see also the discussion in [20]).
@@
Properties (i), (iii) and (vi) imply that every contactomorphism of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ contact isotopic to the identity has at least $n$ translated points with respect to $\alpha_{0}$, and at least $2n$ if either $k$ is even or all the translated points are non-degenerate. We thus recover the corresponding result of [17], but not the optimal bound obtained by the first author in [3], where it is proved that every contactomorphism of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ contact isotopic to the identity has at least $2n$ translated points with respect to $\alpha_{0}$.
@@
Suppose that $k$ is prime. Recall that the cohomological index $\operatorname{ind}(A)$ of a subset $A$ of $L_{k}^{2n-1}(\underline{w})$ is the dimension over $\mathbb{Z}_{k}$ of the image of the map $\check{H}^{\ast}\big{(}L_{k}^{2n-1}(\underline{w});\mathbb{Z}_{k}\big{)}\rightarrow\check{H}^{\ast}(A;\mathbb{Z}_{k})$ on Čech cohomology induced by the inclusion $A\hookrightarrow L_{k}^{2n-1}(\underline{w})$. As we will see, property (iii) can be refined in this case as follows: if
@@
As a first application of Theorem 1.1 we prove that the standard Reeb flow on lens spaces with equal weights is a geodesic for the discriminant and oscillation norms introduced in [10]. The definition of the discriminant norm $\nu_{\operatorname{dis}}$ and of the oscillation pseudonorm $\nu_{\operatorname{osc}}$ on the universal cover of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ are recalled in Section 4 below, as well as the definition of the discriminant and oscillation lengths of contact isotopies. Recall also from [10, Proposition 3.2] that the oscillation pseudonorm is non-degenerate if and only if $(M,\xi)$ is orderable; in particular, it is thus a norm for lens spaces. As in [6], we say that a contact isotopy of a closed orderable contact manifold is a geodesic for the discriminant or for the oscillation norm if its discriminant or oscillation length is equal to the discriminant or oscillation norm of the element of the universal cover it represents. In other words, a contact isotopy is a geodesic for the discriminant or oscillation norm if it minimizes the discriminant or oscillation length in its homotopy class with fixed endpoints. In [6] it is proved that certain contact isotopies of $(\mathbb{R}^{2n}\times\mathbb{S}^{1},\xi_{0})$ are geodesics for the discriminant and oscillation norms. We obtain here a similar result for lens spaces with equal weights, answering a question in [10]. In [10] and [17] respectively it is proved that the discriminant and oscillation norms on real projective spaces and on general lens spaces are unbounded, by showing that the classes in the universal cover represented by higher iterations of the Reeb flow have bigger and bigger discriminant and oscillation norms. More precisely, it is proved in [17] that for every $N$ the discriminant and oscillation norms on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ of $\widetilde{r_{6\pi N}}$ and $\widetilde{r_{20\pi N}}$ respectively are at least equal to $N+1$. Since the discriminant length of $\{r_{6\pi Nt}\}_{t\in[0,1]}$ is $13Nk+$ 1 and the oscillation length of $\{r_{20\pi Nt}\}_{t\in[0,1]}$ is $110Nk+$ 1, the results in [17] (as well as the previous ones in [10]) left open the question of whether there exist contact isotopies in the same homotopy class with fixed endpoints as $\{r_{6\pi Nt}\}_{t\in[0,1]}$ or $\{r_{20\pi Nt}\}_{t\in[0,1]}$ having shorter discriminant or oscillation lengths. Note that this is what happens for the sphere $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$: the $N$-th iteration $\{r_{2\pi Nt}\}_{t\in[0,1]}$ of the Reeb flow $\{r_{2\pi t}\}_{t\in[0,1]}$ of $\bar{\alpha}_{0}$ has discriminant and oscillation length $N+1$, but by [10, Proposition 4.3] the discriminant norm and the oscillation pseudonorm of [$[\{r_{2\pi Nt}\}_{t\in[0,1]}]$ are smaller or equal than $4$; in other words, there exist contact isotopies of $(\mathbb{S}^{2n-1},\bar{\xi}_{0})$ in the same homotopy class with fixed endpoints as certain iterations the Reeb flow having strictly shorter discriminant and oscillation length. As an application of Theorem 1.1, in Section 4 we show that for lens spaces with equal weights this is not possible. More precisely, we show that Theorem 1.1 implies the following result.
@@
For every real number $T$, the Reeb flow $\{r_{Tt}\}_{t\in[0,1]}$ of the standard contact form $\alpha_{0}$ on a lens space of the form $L_{k}^{2n-1}(w,\cdots,w)$ is a geodesic for the discriminant and oscillation norms. In particular,
@@
For general lens spaces $L_{k}^{2n-1}(\underline{w})$ we also obtain lower bounds for the discriminant and oscillation norms of the standard Reeb flow that are sharper than those in [10, 17], however in general these bounds are not sharp enough to prove that the Reeb flow is a geodesic (see Section 4).
Using the spectral selectors of Theorem 1.1 we also define a stably unbounded conjugation invariant pseudonorm on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. More precisely, posing $c_{-}=c_{-2n+1}$ and $c_{+}=c_{0}$ we prove the following result.